OpenStudy (anonymous):

does anyone know this? sqrt-25 / (5-2i)+(1-4i)

OpenStudy (anonymous):

answer choices

OpenStudy (anonymous):

??

OpenStudy (anonymous):

r there answer choices

OpenStudy (anonymous):

the quantity of negative five plus fifteen i all over sixteen the quantity of negative five plus five i all over twelve. the quantity of negative one plus three i all over four. the quantity of five minus five i over twelve

OpenStudy (anonymous):

haha ok

OpenStudy (anonymous):

−25−−−−√(5−2i)+(1−4i) Consider the denominator, (5-2i) + (1-4i) = 6 - 6i Consider the numerator, −25−−−−√=5i So, −25−−−−√(5−2i)+(1−4i) =5i6−6i =5i6(1−i) Multiply the fraction by a conjugate 1+i1+i =5i6(1−i)×1+i1+i =5i(1+i)6(1−i)(1+i) =5i+5i26(12−(−1)) =5i+5(−1)6(2) =5i−512 so this is when it stops lagging i think lol so does this help

OpenStudy (anonymous):

yes it does , Thanks!

OpenStudy (anonymous):

Can u help with one more?

OpenStudy (anonymous):

sure ask away

OpenStudy (anonymous):

What is the 3rd term when -2h4 + 5h2 - 3 - 2h - 8h3 is arranged in descending order? h -8h3 6f4 5h2

OpenStudy (anonymous):

Given f(x) = the quantity of x plus 4, divided by 6, solve for f-1(7). 1 24 32 38

OpenStudy (anonymous):

Can u help with those

OpenStudy (anonymous):

r u in 8 th grade?

OpenStudy (anonymous):

lol no, im on 11th

OpenStudy (anonymous):

remember below :- 1) f^-1 means they want u find x 2) f means they want u find f(x)

OpenStudy (anonymous):

i am in ninth and i learned this last yr

OpenStudy (anonymous):

well im dumb

OpenStudy (anonymous):

lol noooo i am just in a school district that spedd every thing up my sister is doing whut ud call 10 grd worg but here we call it 3rd grade work lol

OpenStudy (anonymous):

ohh ok haha gotcha, do uk how to do the descending ordder one?

OpenStudy (anonymous):

hold on the second one is d

OpenStudy (anonymous):

now let me figure out the second one

OpenStudy (anonymous):

so whut u do is put them in order from hiest to lowest

OpenStudy (anonymous):

go ahead and do that @JohnA8421

OpenStudy (anonymous):

thanks! and ok 12h^2,-8h3, -3, 5h2, -2h4

OpenStudy (anonymous):

check that agian

OpenStudy (anonymous):

were did you get 12h^2

OpenStudy (anonymous):

i meant 2h4, would the answer be B?

OpenStudy (anonymous):

i am gunna repost the question so i dont have to go all the way up to look and all the way down to right

OpenStudy (anonymous):

wright*

OpenStudy (anonymous):

What is the 3rd term when -2h4 + 5h2 - 3 - 2h - 8h3 is arranged in descending order? h -8h3 6f4 5h2

OpenStudy (anonymous):

how did you get b explain yes it is b but explain how you got it

OpenStudy (anonymous):

yes the answer is b

OpenStudy (anonymous):

but explain how u got it

OpenStudy (anonymous):

its b because when its put in order its 5h2, --3,-8h3

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

yes because 5h2 is the only positive so it goes first then id bee -2h4 bc it is the hiest off all negatives then -8h bc it has the variables so that makes it hies than the -3 with no variable so it is b but you did it rong @JohnA8421

OpenStudy (anonymous):

OMG....talk about a tutor

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

need mor help @JohnA8421

OpenStudy (anonymous):

hmm, yea i might need more help in a lil bit

OpenStudy (anonymous):

ok look up there^ i replied to ur last answer

OpenStudy (anonymous):

i saw it, thanks!

OpenStudy (anonymous):

yeparoo